• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 21:03:09 on June 12, 2009

    Entry number 280864

    keyword=Compton electron detector access and moved in Compton position

    We added back the additionnal inhibit cables, it seems a group of input might have a hardware problem. So Inhibit signal is on input 4,5 and 6. Also the rates seemed to come from the pulser, it seems there is a bug in the controls to turn it off, will double check at next access. We physically unplugged the signal. The DAQ should not crash anymore and the high rate will hopefully disappear.